OpenStudy (anonymous):

What is 2s-t if s=9 and t=12 A)-6 B) 6 C) 15 D) 30

OpenStudy (colcaps):

B 2(9) - 12 Apply PEMDAS so 9 x 2 = 18 18 -12 = 6 ans. B

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What is a-24 if a=6 A)-30 B) -18 C) 18 D)22 I got b but idk

OpenStudy (askmelots):

6-24=-18 you are correct

OpenStudy (anonymous):

Thank you could you help me once or twice more

OpenStudy (anonymous):

@Askmelots

OpenStudy (askmelots):

yes, sorry I was helping others